OpenStudy (anonymous):

Find and classify the critical point(s) f(x) = x^3 - x^2 - x I know I need to take the derivative which would give me 3x^2 - 2x - 1 = 0 and I think I need to set them to zero to find the points but I'm not sure? would it be like 3x^2 = 0 and 2x-1 = 0 and then solve?

OpenStudy (samgrace):

Critical points are where the graph intersects the dependent axis, the x axis. Therefore derivatives are not the correct course of action, these would give you the minimum/maximum points.

OpenStudy (samgrace):

Put the graph equal to zero and solve for x by factoring out x

OpenStudy (samgrace):

\[f(x) = x^3 - x^2 - x => 0 = x(x^2 - x - 1) \] \[x = (-b \pm \sqrt(b^2-4ac))/2a\]

OpenStudy (anonymous):

wait, what numbers do I plug in for a, b, c then? 3, 2, 1?

OpenStudy (anonymous):

Take the derivative as you did and then factor the derivative. 3x^2 - 2x - 1 = 0 is just a regular quadratic that can be factored. If you factor it correctly, the values for x that you find that makes the derivative equal 0 will be the x-coordinates of the critical points youre looking for.

OpenStudy (anonymous):

i think the critical points are where the derivative is equal to zero, in the case of a polynomial take the derivative, set it equal to zero and solve for \(x\)

OpenStudy (anonymous):

if you cannot see how to factor it, or if it does not factor, use the quadratic formula i suck at factoring, so i almost always use it